# The Scaling Law in Stellar Light Curves

### v1

@@ 1 Introduction
Despite the flood of time series data in astronomy, characterizing these light curves to perform downstream label inferences or identify out-of-distribution objects remains an unresolved problem. Traditional methods often rely on crude summary statistics (Bastien et al., 2013). For example, in the field of asteroseismology, which involves characterizing light curves from stars, the established method calculates the power spectrum of the light curves and extracts the frequency of the maximum oscillation power, $\nu_{\rm max}$ , which highly correlate with stellar masses and ages (Aerts et al., 2010). Therefore, asteroseismology has been the primary way to achieve golden measurements of how old stars are. However, such summary statistics do not fully harness all the information contained in the light curves and are often difficult to determine due to heterogeneous noise levels (Stello et al., 2022) and irregular cadences (Hambleton et al., 2023).
@@
To overcome this limitation, a growing body of literature has developed machine learning methods for time-domain astronomy (Blancato et al., 2022; Pan et al., 2024). In particular, for stellar light curves, Transformer models (Vaswani et al., 2023) is receiving an increasing interest (Donoso-Oliva et al., 2023). For example, Pan et al. (2024) customized Transformer models for light curves and demonstrated that, in the case of extracting the surface gravity (or $\log g$) of stars, the ability of Transformer models to extract long-range information can yield state-of-the-art performance in a supervised learning setting.
@@
While deep learning has shown promise in characterizing stellar light curves, most studies have been limited to discriminative supervised learning (Hon et al., 2017; Sayeed et al., 2021). Such methods often fall short in applications due to two main reasons. First, there is a lack of existing ground truth labels in the data. For example, while the Rubin Observatory will collect on the order of 10 billion time series data, the number of existing labels (obtained from other more observationally expensive means – such as monitoring eclipsing binaries) will remain on the order of thousands (Kirk et al., 2016). Second, there is a need for a generative model that can extract robust, generalizable representations suitable for other downstream tasks and out-of-distribution search. And solving these bottlenecks is critical to limit redundant resource dedications for different pipelines (Hon et al., 2017; Sayeed et al., 2021; Pan et al., 2024) and downstream tasks.
@@
This limitation has called for a different approach that can utilize unlabeled data and can generalise for various downstream tasks. A prominent candidate would be training foundational models through self-supervised learning as such models have shown promises in different domains such as text data and images (Brown et al., 2020; Bai et al., 2023b). These models, especially Large Language Models (LLMs), exhibit performance unmatched by their smaller variants and emergent abilities such as in-context learning (Brown et al., 2020). Furthermore, in recent years, foundational models in other domains has observed the scaling law, which is the empirical observation that the next-token prediction loss of Transformer-based autoregressive generative models scales in a power law with increasing compute budgets, model size and the number of training tokens (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022). Applying the same idea to astronomical time series data invites two questions: (1) Would the same self-supervised generative models based on Transformer exhibit the scaling law on astronomical time series data? (2) Would a better next-token prediction loss obtained by scaling up translate to better performance in interested downstream tasks?
@@ 2 Kepler Mission’s Stellar Light Curves
In this study, we focus on the stellar light curves from the Kepler mission. The Kepler mission was a flagship NASA mission that operated from 2009 to 2013. It yielded 4 years of light curves for approximately 200,000 stars at an equal cadence of about 29.4 minutes, containing $\sim$70,000 timestamps and corresponding observed fluxes. In order to compare with the supervised deep learning Transformer-based models of (Pan et al., 2024), we follow the same Kepler data selection, selecting 17,201 high-quality light curves as our dataset. Some light curve examples are shown in Fig 1. The different panels show the cases with four representative stars with different surface gravity values, where a smaller $\log g$ indicates a larger stellar radius and, consequently, less oscillation in the stellar fluxes.
The light curves are selected based on the availability of the asteroseismic estimated $\log g$ values for these stars, as provided in (Mathur et al., 2017; Yu et al., 2018, 2020). These studies focus on extracting the properties of stars with the maximum oscillation frequency $\nu_{\rm max}$, and the prominence of $\nu_{\rm max}$ makes it an effective way to select light curves that have high signal-to-noise ratios. Furthermore, the radii of the stars are also provided in Berger et al. (2020), which allow for the application of a radius-dependent high-pass filter to remove instrumental effects (García et al., 2011).
@@ 3 Learning Stellar Light Curves with GPT-2
GPT-2 is a decoder-only Transformer with learnable and straightforward positional embeddings and Pre-layer Normalization, in contrast to more recent LLM architecture such as Llama (Touvron et al., 2023) and Qwen (Bai et al., 2023a), which instead use Rotary Positional Embedding (Su et al., 2023) and RMSNorm (Zhang & Sennrich, 2019). We train all our models following the training scheme in (Radford et al., 2019), including the AdamW optimizer (Loshchilov & Hutter, 2019). The learning rate is increased linearly to the peak value in 2,000 iterations and then annealed to $2\times 10^{-4}$ with a cosine schedule while ensuring the learning rate remains higher than $2\times 10^{-5}$. We also adopt the default batch size of 12 and accumulate gradients for 40 batches, leading to an effective batch size of 480 light curve patches or 38,400 tokens. All models are trained for 200,000 iterations, summing up to 7.7B processed tokens.
@@
In addition to the original GPT-2, GPT-2 medium, GPT-2 large, and GPT-2 XL models proposed in (Radford et al., 2019), we also train models of different sizes by adjusting the number of layers (depth), number of self-attention heads (Head), and hidden dimension size (width), as listed in Table 1. This is to ensure that we can evaluate the performance of the models with a larger dynamic range in terms of model size, ranging from $10^{4}$ to $10^{9}$ parameters. We note that the largest model that we train, with $\sim 1.5$B parameters, starts to rival some of the “lightweight” modern-day LLMs (Bai et al., 2023a).
@@ 4 Results
The study of Kepler’s light curves is a cornerstone in modern-day stellar astrophysics. The temporal variations of the fluxes of stars are the manifestation of physical processes where light bounces within the stellar interior. As such, the light curves are tell-tale signatures of the stellar interior physics that are otherwise inaccessible to us. In this study, we train GPT-2-like models as autoregressive generative models to capture this inherent interior physics. Note that, the models are all initialized completely randomly from $\mathcal{N}(0,0.02)$ to study the emergence of the scaling law in the generative models training from scratch.
We will start by showcasing the qualitative results from the GPT-2 XL model with $1.5$B parameters. In Figure 1, we show the one-step predictions of all fluxes, where at each step, we perform the prediction conditioning on all the previous (true) observed fluxes. We note that, given the predominantly stochastic nature of stellar flux variations, i.e., given all $N$ previous steps, the next “token” prediction complies to a distribution, just like in natural languages. As such, we do not expect the predictions to be exact. Specifically, since our prediction is based upon maximal likelihood estimation from a multi-mode distribution, the multi-step prediction would deviate from the truly observed fluxes although the predicted light curve could be statistical the same as the original light curves. Due to this consideration, we only examine the trend with one-step predictions.
As shown in Figure 1, like in natural languages, the autoregressive generative model does learn the “stylistic patterns” of the stellar light curves, where it shows a good understanding of the long timescale oscillations from stars with small surface gravity as opposed to short oscillations from stars with large surface gravity. More importantly, the ability to generally grasp the patterns of the light curves suggests that we could further improve the predictions if a scaling law applies to the model, and the learned latent representation might be robust for further downstream tasks such as predicting the surface gravity (and hence ages) of the stars.
Figure 2: The emergence of scaling law for stellar light curves. We train autoregressive generative models on stellar light curves with different model sizes, ranging from $10^{4}$ to $10^{9}$ parameters. The different lines show the pretraining loss (Huber loss of next-token prediction) as a function of the computational cost for different model sizes. We truncate the training of the models when the slope of the tangent line connecting the loss from the last three epochs becomes shallower than the overall scaling law. The prediction loss plateaus at lower values for larger models, demonstrating that the scaling law also applies to Transformer-based autoregressive generative models when applied to stellar light curves.
4.1 The Emergent of Scaling Law for Light Curves
@@
We investigate the scaling law between pretraining loss and compute budget with model sizes spanning $10^{4}-10^{9}$ parameters. Figure 2 shows the pretraining loss of the test set as a function of the computation budget in PetaFlop(PF)-days2221 PF-day = $10^{15}\times 24\times 3600=8.64\times 10^{19}$ floating point operations $\approx 3.2$ A100 days. for different GPT-2-like models, with individual lines of different colors representing the test loss curve for various computational budgets. We define the critical point as 3.4 epoch, where the local slope of the loss curve becomes shallower than the overall scaling law defined below. The models are trained for 11.4 epochs. The loss curves before the critical point are shown in solid colors, while the curves after the critical points are depicted with more translucent colors.
Following (Henighan et al., 2020), we fit a power law to the critical points of of the loss as a function [^3] The original expression in (Henighan et al., 2020) is $L(C)=(\frac{C_{0}}{C})^{\alpha_{C}}$, we apply a base-10 logarithm to the expression for better readability. Throughout this study, we consistently utilize the modified equivalent expression. of the computational cost $C$, $\log_{10}L(C)=\alpha_{C}\log_{10}C+\beta_{C}$ with $\alpha_{C}=-0.00448,\beta_{C}=-2.386$, demonstrating the emergence of a scaling law in stellar light curve data. The scaling suggests that, for a 10 fold larger in compute budgets, the next-token prediction Huber loss (again, a rectified version of MSE) would improve by 1%. Already, we can predict that next-token prediction Huber loss will improve 3.8% compared to our most capable 1.5B GPT-2 XL model at a compute budget of $10^{4}$ PF-days.
We suspect the somewhat shallow slope (i.e., smaller gain as model size increases) is due to the limited training size, as scaling laws observed in (Kaplan et al., 2020) are also functions of the dataset size. We aim to explore the dependency on the training set size in our future work.
@@ 4.2 Scaling Law in Inferring Stellar Surface Gravity
For the label inference head (mapping from the last layer representation to $\log g$), we adopt a 3-layer MLP with a hidden dimension of 64. As different models have different width , the dimensions for the representation and hence the input dimension of the MLP head also vary. The largest MLP head has 0.11M parameters. We freeze the main generative model and train using a dataset consisting of 10,000 patches for training and 20,000 patches for validation. As mentioned in Section 2, our pretraining dataset contains over 8 million patches from more than 16,000 stars. The 100,00 patches used for the downstream task are selected from this pretraining dataset, corresponding to about $10,000$ star. This sample size is representative of the number of patches that could realistically have golden labels obtained through other means in a practical setting. To ensure no information leakage, we use a separate test set of 20,000 patches drawn from the test set, which was not used during the pretraining of the generative models.
@@
We compare our results from self-supervised learning with GPT-2 models to the current state-of-the-art (SOTA) supervised learning method, Astroconformer (Pan et al., 2024), represented by the solid orange horizontal lines. Astroconformer is a Transformer-based model with 0.86M parameters that learns from the direct relationship between the stellar light curve and $\log g$. They argue that introducing a strong inductive bias is necessary because the labeled training sample is limited. To address this, they combine convolution and self-attention mechanisms to simultaneously capture localized and global correlations, which they found to outperform other existing supervised learning methods primarily based on convolutional neural networks or K-nearest neighbor search (Sayeed et al., 2021). To ensure a fair comparison, the Astroconformer results presented here are trained with the same labeled training samples used for evaluating self-supervised representations. Further, to explore the sample efficiency gain by self-supervised learning, we also evaluate Astroconformer’s performance based on 10 times more training samples drawn from the pretraining dataset.
@@
The figure shows that the scaling law enables sufficiently large self-supervised autoregressive models, even with the 0.7B pre-training tokens, to surpass the performance of the current supervised learning models. Astroconformer reaches a $\log g$ prediction MSE of 0.083 and 0.059 when trained on 10,000 samples and 100,000 samples, respectively. Even with the 85M parameters GPT-2 model, the learned latent representation is robust to surpass supervised learning trained on 10 times of data, which indicates they are $>10$ more label-efficient. , We fit the power law as before, and find the scaling law for the $\log g$ prediction MSE follows $\log_{10}L(N)=-0.063\log_{10}\log N-0.725$. In other words, if we extrapolate that to the typical industrial scale of 10B or 100B, the $\log g$ precision MSE would be 11% and 23% lower than our most capable 1.5B GPT-2 XL model.
@@
Figure 5: The impact of the labeled training dataset on the downstream $\log g$ inference task is illustrated in Figure 5. We compare the performance of GPT-2 XL with Astroconformer. For any given fixed set of labeled data, the autoregressive generative approach adopted in this study leads to better precision than the supervised method. The self-supervised learning approach requires less than one third of training samples to achieve the same precision in $\log g$ inference.

### v2

@@ 1 Introduction
Despite the flood of time series data in astronomy, characterizing these light curves to perform downstream label inferences or identify out-of-distribution objects remains an unresolved problem. Traditional methods often rely on crude summary statistics (Bastien et al., 2013). For example, in the field of asteroseismology, which involves characterizing light curves from stars, the established method calculates the power spectrum of the light curves and extracts the frequency of the maximum oscillation power, $\nu_{\rm max}$ , which highly correlates with stellar masses and ages (Aerts et al., 2010). Therefore, asteroseismology has been the primary way to achieve golden measurements of how old stars are. However, such summary statistics do not fully harness all the information contained in the light curves and are often difficult to determine due to heterogeneous noise levels (Stello et al., 2022) and irregular cadences (Hambleton et al., 2023).
@@
To overcome this limitation, a growing body of literature has developed machine learning methods for time-domain astronomy (Blancato et al., 2022; Pan et al., 2024). In particular, for stellar light curves, Transformer models (Vaswani et al., 2017) is receiving an increasing interest (Donoso-Oliva et al., 2023). For example, Pan et al. (2024) customized Transformer models for light curves and demonstrated that, in the case of extracting the surface gravity (or $\log g$) of stars, the ability of Transformer models to extract long-range information can yield state-of-the-art performance in a supervised learning setting.
@@
While deep learning has shown promise in characterizing stellar light curves, most studies have been limited to discriminative supervised learning (Hon et al., 2017; Sayeed et al., 2021). Such methods often fall short in applications due to two main reasons. First, there is a lack of existing ground truth labels for the data. For example, while the Rubin Observatory will collect tens of billions of time series data, the number of existing labels (obtained from other more observationally expensive means – such as monitoring eclipsing binaries) will remain on the order of thousands (Kirk et al., 2016). Second, there is a need for a generative model that can extract robust, generalizable representations suitable for other downstream tasks and out-of-distribution search. And solving these bottlenecks is critical to limit redundant resource dedications for different pipelines (Hon et al., 2017; Sayeed et al., 2021; Pan et al., 2024) and downstream tasks.
@@
This limitation has called for a different approach that can utilize unlabeled data and can generalise for various downstream tasks. A prominent candidate would be training foundational models through self-supervised learning as such models have shown promises in different domains such as text data and images (Brown et al., 2020; Bai et al., 2023b). These models, especially Large Language Models (LLMs), exhibit performance unmatched by their smaller variants and emergent abilities such as in-context learning (Brown et al., 2020). Furthermore, in recent years, foundational models in other domains has observed the scaling law, which is the empirical observation that the next-token prediction loss of Transformer-based autoregressive generative models scales in a power law with increasing compute budgets, model size and the number of training tokens (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022). Applying the same idea to astronomical time series data invites two questions: (1) Would the same self-supervised generative models based on Transformer exhibit the scaling law on astronomical time series data? (2) Would a better next-token prediction loss obtained by scaling up translate to better performance in downstream tasks of interest?
@@ 2 Kepler Mission’s Stellar Light Curves
In this study, we focus on the stellar light curves from the Kepler mission. The Kepler mission was a flagship NASA mission that operated from 2009 to 2013. It yielded 4 years of light curves for approximately 200,000 stars at an equal cadence of about 29.4 minutes, containing $\sim$70,000 timestamps and corresponding observed fluxes. In order to compare with the supervised deep learning Transformer-based models (Pan et al., 2024), we follow the same Kepler data selection, selecting 17,201 high-quality light curves as our dataset. Some light curve examples are shown in Fig 1. The different panels show the cases with four representative stars with different surface gravity values, where a smaller $\log g$ indicates a larger stellar radius and, consequently, less oscillation in the stellar fluxes.
The light curves are selected based on the availability of the asteroseismic $\log g$ values for these stars, as provided in previous studies (Mathur et al., 2017; Yu et al., 2018, 2020). These studies focus on extracting the properties of stars with the maximum oscillation frequency $\nu_{\rm max}$, and the prominence of $\nu_{\rm max}$ makes it an effective way to select light curves that have high signal-to-noise ratios. Furthermore, the radii of the stars are also provided in Berger et al. (2020), which allow for the application of a radius-dependent high-pass filter to remove instrumental effects (García et al., 2011).
@@ 3 Learning Stellar Light Curves with GPT-2
GPT-2 is a decoder-only Transformer with learnable and straightforward positional embeddings and Pre-layer Normalization, in contrast to more recent LLM architecture such as Llama (Touvron et al., 2023) and Qwen (Bai et al., 2023a), which instead use Rotary Positional Embedding (Su et al., 2023) and RMSNorm (Zhang & Sennrich, 2019). We train all our models following the training scheme in Radford et al. (2019). We adopt the AdamW optimizer (Loshchilov & Hutter, 2019). The learning rate is increased linearly to the peak value in 2,000 iterations and then annealed to $2\times 10^{-4}$ with a cosine schedule while ensuring the learning rate remains higher than $2\times 10^{-5}$. We also adopt the default batch size of 12 and accumulate gradients for 40 batches, leading to an effective batch size of 480 light curve patches or 38,400 tokens. All models are trained for 200,000 iterations, summing up to 7.7B processed tokens.
@@
In addition to the original GPT-2, GPT-2 medium, GPT-2 large, and GPT-2 XL models proposed in Radford et al. (2019), we also train models of different sizes by adjusting the number of layers (depth), number of self-attention heads (Head), and hidden dimension size (width), as listed in Table 1. This is to ensure that we can evaluate the performance of the models with a larger dynamic range in terms of model size, ranging from $10^{4}$ to $10^{9}$ parameters. We note that the largest model that we train, with $\sim 1.5$B parameters, starts to rival some of the “lightweight” modern-day LLMs (Bai et al., 2023a).
@@ 4 Results
The study of Kepler’s light curves is a cornerstone in modern-day stellar astrophysics. The temporal variations of the fluxes of stars are the manifestation of physical processes where light bounces within the stellar interior. As such, the light curves are tell-tale signatures of the stellar interior physics that are otherwise inaccessible to us. In this study, we train GPT-2-like models as autoregressive generative models to capture this inherent interior physics. Note that, the models are all initialized completely randomly from $\mathcal{N}(0,0.02)$ to study the emergence of the scaling law in the generative models trained from scratch.
We will start by showcasing the qualitative results from the GPT-2 XL model with $1.5$B parameters. In Figure 1, we show the one-step predictions of all fluxes, where at each step, we perform the prediction conditioning on all the previous (true) observed fluxes. We note that, given the predominantly stochastic nature of stellar flux variations, i.e., given all $N$ previous steps, the next “token” prediction complies to a distribution, just like in natural languages. As such, we do not expect the predictions to be exact. Specifically, since our prediction is based upon maximal likelihood estimation from a multi-mode distribution, the multi-step prediction would deviate from the truly observed fluxes although the predicted light curve could be statistically the same as the original light curves. Due to this consideration, we only examine the trend with one-step predictions.
As shown in Figure 1, like in natural languages, the autoregressive generative model does learn the “stylistic patterns” of the stellar light curves, where it shows a good understanding of the long timescale oscillations from stars with small surface gravity as opposed to short oscillations from stars with large surface gravity. More importantly, the ability to generally grasp the patterns of the light curves suggests that we could further improve the predictions if a scaling law applies to the model, and the learned latent representations might be robust for further downstream tasks such as predicting the surface gravity (and hence ages) of the stars.
Figure 2: The emergence of scaling law for stellar light curves. We train autoregressive generative models on stellar light curves with different model sizes, ranging from $10^{4}$ to $10^{9}$ parameters. The different lines show the pretraining loss (Huber loss of next-token prediction) as a function of the compute budget for different model sizes. We truncate the training of the models when the slope of the tangent line connecting the loss from the last three epochs becomes shallower than the overall scaling law. The prediction loss plateaus at lower values for larger models, demonstrating that the scaling law also applies to Transformer-based autoregressive generative models when applied to stellar light curves.
4.1 The Emergence of Scaling Law for Light Curves
@@
We investigate the scaling law between pretraining loss and compute budget with model sizes spanning $10^{4}-10^{9}$ parameters. Figure 2 shows the pretraining loss of the test set as a function of the compute budget in PetaFlop(PF)-days2221 PF-day = $10^{15}\times 24\times 3600=8.64\times 10^{19}$ floating point operations $\approx 3.2$ A100 days. for different GPT-2-like models, with individual lines of different colors representing the test loss curve for various compute budgets. We define the critical point as 3.4 epoch, where the local slope of the loss curve becomes shallower than the overall scaling law defined below. The models are trained for 11.4 epochs. The loss curves before the critical point are shown in solid colors, while the curves after the critical points are depicted with more translucent colors.
Following Henighan et al. (2020), we fit a power law to the loss at the critical points as a function [^3] The original expression in Henighan et al. (2020) is $L(C)=(\frac{C_{0}}{C})^{\alpha_{C}}$, we apply a base-10 logarithm to the expression for better readability. Throughout this study, we consistently utilize the modified equivalent expression. of the compute budget $C$, $\log_{10}L(C)=\alpha_{C}\log_{10}C+\beta_{C}$ with $\alpha_{C}=-0.00448,\beta_{C}=-2.386$, demonstrating the emergence of a scaling law in stellar light curve data. The scaling suggests that, for a 10 fold larger compute budget, the next-token prediction Huber loss (again, a rectified version of MSE) would improve by 1%. Already, we can predict that next-token prediction Huber loss will improve 3.8% compared to our most capable 1.5B GPT-2 XL model at a compute budget of $10^{4}$ PF-days.
We suspect the somewhat shallow slope, i.e., smaller gain as model size increases, is due to the limited training data size, as scaling laws observed in (Kaplan et al., 2020) are also functions of the dataset size. We aim to explore the dependency on the training set size in our future work.
@@ 4.2 Scaling Law in Inferring Stellar Surface Gravity
For the label inference head (mapping from the last layer representation to $\log g$), we adopt a 3-layer MLP with a hidden dimension of 64. As different models have different width, the dimensions for the representation and hence the input dimension of the MLP head also varies. The largest MLP head has 0.11M parameters. We freeze the main generative model and train using a dataset consisting of 10,000 patches for training and 20,000 patches for validation. As mentioned in Section 2, our pretraining dataset contains over 8 million patches from more than 16,000 stars. The 10,000 patches used for the downstream task are selected from this pretraining dataset, corresponding to about $10,000$ stars. This sample size is representative of the number of golden labels obtained through other means in a practical setting. To ensure no information leakage, we use a separate test set of 20,000 patches drawn from the test set, which was not used during the pretraining of the generative models.
@@
We compare our results from self-supervised learning with GPT-2 models to the current state-of-the-art (SOTA) supervised learning method, Astroconformer (Pan et al., 2024), represented by the solid orange horizontal lines. Astroconformer is a Transformer-based model with 0.86M parameters that learns from the direct relationship between the stellar light curves and $\log g$. They argue that introducing a strong inductive bias is necessary because the labeled training sample is limited. To address this, they combine convolution and self-attention mechanisms to simultaneously capture localized and global correlations, which they found to outperform other existing supervised learning methods primarily based on convolutional neural networks or K-nearest neighbor search (Sayeed et al., 2021). To ensure a fair comparison, the Astroconformer results presented here are trained with the same labeled training samples used for evaluating self-supervised representations. Further, to explore the sample efficiency gain by self-supervised learning, we also evaluate Astroconformer’s performance based on 10 times more training samples drawn from the pretraining dataset.
@@
The figure shows that the scaling law enables sufficiently large self-supervised autoregressive models, even with the 0.7B pretraining tokens, to surpass the performance of the current supervised learning models. Astroconformer reaches a $\log g$ prediction MSE of 0.083 and 0.059 when trained on 10,000 samples and 100,000 samples, respectively. Even with the 85M parameters GPT-2 model, the learned latent representation is robust to surpass supervised learning trained on 10 times of data, which indicates they are $>10$ times more label-efficient. We fit the power law as before, and find the scaling law for the $\log g$ prediction MSE follows $\log_{10}L(N)=-0.063\log_{10}N-0.725$. In other words, if we extrapolate that to the typical industrial scale of 10B or 100B, the $\log g$ precision MSE would be 11% and 23% lower than our most capable 1.5B GPT-2 XL model.
@@
Figure 5: The impact of the labeled training dataset on the downstream $\log g$ inference task. We compare the performance of GPT-2 XL with that of Astroconformer. For any given fixed set of labeled data, the autoregressive generative approach adopted in this study leads to better precision than the supervised method. The self-supervised learning approach requires less than one third of training samples to achieve the same precision in $\log g$ inference.